### v1

# MMDesign: Multi-Modality Transfer Learning for Generative Protein Design

Abstract
Protein design involves generating protein sequences based on their corresponding protein backbones. While deep generative models show promise for learning protein design directly from data, the lack of publicly available structure-sequence pairings limits their generalization capabilities. Previous efforts of generative protein design have focused on architectural improvements and pseudo-data augmentation to overcome this bottleneck. To further address this challenge, we propose a novel protein design paradigm called MMDesign, which leverages multi-modality transfer learning. To our knowledge, MMDesign is the first framework that combines a pretrained structural module with a pretrained contextual module, using an auto-encoder (AE) based language model to incorporate prior semantic knowledge of protein sequences. We also introduce a cross-layer cross-modal alignment algorithm to enable the structural module to learn long-term temporal information and ensure consistency between structural and contextual modalities. Experimental results, only training with the small CATH dataset, demonstrate that our MMDesign framework consistently outperforms other baselines on various public test sets. To further assess the biological plausibility of the generated protein sequences and data distribution, we present systematic quantitative analysis techniques that provide interpretability and reveal more about the laws of protein design.
Figure 1: (a) The mainstream advanced deep protein design framework, where the GNN-based structure module represents the structural features, followed by an optional contextual module encoding the sequence information, and the decoder responsible for decoding the generated protein sequences. (b) The proposed novel paradigm of the protein design framework consists of two pretrained modules, i.e., the pretrained graph structure module and the pretrained auto-encoder (AE) contextual module. Additional cross-layer cross-modal alignment aims to enhance feature constraints.
1 Introduction
De novo protein design, also known as protein inverse folding, aims to generate protein sequences (i.e., amino acids) based on the corresponding atomic coordinates of the protein backbones. Protein design is crucial for various applications in bioengineering, including drug design and enzyme discovery, but it remains one of the challenging tasks. Recent years have seen significant progress in protein design, thanks to the emergence of deep learning-based methods. In particular, supervised generative models have enabled more efficient and effective prediction of protein sequences[13, 23, 2, 14, 8, 4]. Deep generative models have the potential to learn protein design rules directly from data, making them a promising alternative to traditional approaches such as physics-based energy functions.
The current mainstream and advanced generative protein design frameworks typically follow an encoder-decoder architecture, as illustrated in Figure 1(a). Firstly, a graph is used to represent the structural information in Euclidean space based on the atomic coordinates of the protein backbone. Subsequently, graph neural networks (GNNs) are responsible for encoding graph-based geometric information into higher dimensional representations using a message-passing mechanism. Finally, the decoder module generates the protein sequences through either autoregressive networks based on self-attention[13, 14] or non-autoregressive networks based on MLPs[5, 6, 32]. However, the lack of massively available parallel datasets of protein structure-sequence pairings has become a major bottleneck for deep protein design models, leading to problems such as under-training or over-fitting. Although billions of protein sequences are available, only a few hundred thousand 3D protein structures have been experimentally determined due to the extremely high costs. Hence, as a multi-modality generative task, it is still challenging to train a practical protein design model based on available low-resource data, especially when compared to other established cross-modal tasks such as vision-language tasks.
To address this limitation, various solutions have emerged: 1) The most common method is to focus on protein design architecture, mainly by enhancing the GNN-based structural module’s encoding capability of protein 3D spatial structure (e.g., introducing a robust geometric constraint network for enhanced equivariance and invariance[13, 14]), or extending the structural module with a temporal module (e.g., RNN or self-attention layers) to further learn relationships between nodes [8], as shown in Figure 1(a). Although only a slight performance improvement, configurations such as equivalent GNNs and temporal modules that source this method have become standard in mainstream protein design models. 2) The most straightforward approach is to expand the dataset by generating large-scale pseudo-structures[8] using advanced protein structure prediction models[15, 3, 18]. Despite data augmentation being widely employed, the data construction process is tedious and time-consuming, and pseudo-data might also introduce potential noise. 3) Another newly emerging framework is to introduce a pretrained structural module, which greatly reduces the reliance on supervised data while maintaining satisfactory performance [32]. Although transfer learning is common in fields such as natural language processing and computer vision, it is not widely used for protein design due to the complexity of structure model pretraining and the limited publicly available related works.
Among these solutions, utilizing pretrained structural networks has become a consensus to overcome the low-resource limit. To further enhance the temporal module that complements the structural module with semantic and temporal representation, it is worth considering injecting the prior language knowledge of protein sequences. However, incorporating prior knowledge of both source and target modalities is intractable for supervised tasks, particularly in protein design where there are no reference works available. Fortunately, a new paradigm in the visual-textual domain [33] allows us to introduce prior knowledge of the target modality in an auto-encoder (AE) manner.
Figure 2: Diagram of the MMDesign framework. The proposed training pipeline is divided into two steps: Step 1 involves pretraining the structural module and the contextual module separately, while Step 2 entails the pretrained modules transferred from Step 1 to optimize the overall MMDesign framework.
Based on the observations above, we propose a novel paradigm for generative protein design, MMDesign, which leverages multi-modality transfer learning through pretrained structural and contextual modules, as illustrated in Figure 1(b). This framework fully exploits pretraining models to overcome the data bottleneck. For the structural module, we incorporate advanced equivariant network-based pretrained structural representation capabilities. For the contextual module, we replace the previous temporal module and decoder module with an AE model, which can integrate comprehensive prior language knowledge and assume the functions of protein semantics encoding and sequence generation.
Furthermore, we incorporate explicit cross-modal alignment constraints to MMDesign, which have been proven effective in multi-modality domains. For example, [20, 7] employ the distillation loss to align the visual and textual modules. This cross-modal alignment can be viewed as a “consistency” metric to measure the distance between two probability distributions [36, 33]. Maintaining consistency across modalities allows the structural module to be constrained by long-range temporal context, as shown in Figure 1(b). Notably, the AE model serves as an implicit cross-modal bridge between the structural and contextual modality, similar to the impact of “consistency”[33].
MMDesign is trained solely on the small dataset CATH [22, 13], yet extensive experiments demonstrate that MMDesign substantially outperforms the state-of-the-art baselines without relying on large-scale supervised data. However, previous studies have focused primarily on average performance, with little attention paid to explaining the generated results and their biological plausibility. We present the first systematic quantitative and comprehensive analysis techniques to assess the plausibility of the generated protein sequences and the distribution of data. The analysis techniques introduced are new for protein design studies, validating the model’s generalization on unseen and unknown datasets, examining the effect of sequence length on the model, assessing the biological plausibility of the generated data distribution and foldability, and performing internal crucial segment discovery and confusion matrix analysis.
The main contributions are as follows:
  • The proposed MMDesign, a novel protein design framework that includes both pretrained structural knowledge and prior contextual semantic knowledge for protein design. This provides a new way of thinking to overcome the bottleneck of other cross-modal protein tasks.
  • The introduction of explicit cross-modal alignment between the structural and contextual modality using a cross-layer consistency constraint and implicit cross-modal alignment using favourable properties of the AE mechanism.
  • Despite only training on a small dataset, MMDesign substantially outperforms other baselines on multiple benchmarks and even surpasses the strong baselines trained on large datasets.
  • More comprehensive interpretable analyses are presented, which provide multidimensional validation and reveal the underlying protein design patterns.
  • The code and trained models will be publicly released.
2 Related Work
The research of protein engineering based on deep learning has attracted more and more attention [12, 11, 10, 19, 9]. Early work on protein design primarily relied on using an energy function that directly models the physical basis of the protein folding state [1]. In recent years, deep learning-based approaches have gained popularity, and generative models that predict the sequences of structures have gradually replaced traditional methods, yielding superior results. For instance, Structured Transformer[13] introduced the formulation of fixed-backbone design as a conditional sequence generation problem. It used invariant features with GNNs and modeled each amino acid as a graph node with edges connecting spatially adjacent amino acids. GVP-GNN [14] further improved GNNs by developing architectures with translation- and rotation-equivariance to enable geometric reasoning, highlighting the powerful structural representation of the geometric vector perceptron (GVP) and GNN module. More recently, ProteinMPNN [4] achieved higher-level sequence recovery on the native backbone by incorporating noise during training and it also improved inference efficiency by performing large-scale side-chain packing calculations. Then PiFold [6] contained a novel residue featurizer and GNN layers to predict protein sequences in a one-shot way. While PiFold improves efficiency and optimizes recovery in single and short chains, it struggles with longer sequences. ESM-IF[8] utilized GVP-based encoder layers to extract geometric features followed by a Transformer. It also employed a pseudo-data augmentation strategy that leveraged the structures of 12 million protein sequences predicted by AlphaFold2[15], achieving an impressive native sequence recovery. However, scaling up the training data by nearly three orders of magnitude significantly increases the training difficulty. Moreover, its better performance does not extend to single and short chains. Overall, the existing works still exhibit limited generalization capabilities without relying on large-scale data. Furthermore, they do not offer a comprehensive biological plausibility analysis of the predicted protein sequences and lack interpretability.
3 Methods
[…]
The overall training pipeline is divided into two main steps. Step 1 is to initialize the GVPConv-based structural module and pretrain the AE network for the contextual module. The parameters of the structural module are directly derived from the off-the-shelf protein structure model [32], which is trained in a self-supervised manner. The contextual module pertaining can be regarded as a sequence-to-sequence pseudo-translation task. And step 2 is to transfer the pretrained structural module and contextual module into the MMDesign framework. To encourage multi-modality feature interaction, we also introduce cross-layer cross-modal constraints.
3.1 Pretraining Stage
3.1.1 Geometric Structural Module Initialization
The GVPConv structural module is a key component of our MMDesign framework. GVP has been widely used in protein design[14, 8] due to its lightweight and efficient linear transformation, as well as its ability to learn rotation-equivariant transformations of vector features and rotation-invariant transformations of scalar features. In our framework, we use a variant GVP as in ESM-IF [8], which transforms the vector features into local reference frames defined for each amino acid to derive rotation-invariant features.
To ensure that the predicted sequence is independent of the reference frame of the structural coordinates, we enforce invariance under any rotation and translation $T$ of the input coordinates $X$, i.e., $p(Y|X)=p(Y|TX)$, where $Y$ is the predicted sequence. The full graph convolution following GVP utilizes message passing to update the node embeddings in each propagation step, taking into account messages from the neighbor nodes and edges [8, 14].
To obtain a better representation of the protein structures, we initialize the GVPConv parameters as [32], where the GVPConv module is pretrained with guidance from a large-scale protein language model through contrastive alignment constraints.
3.1.2 Contextual Transformer Pretraining
Our AE model is a generic auto-regressive encoder-decoder Transformer architecture[27]. The only modification is the use of learned positional embeddings instead of classical sinusoidal positional embeddings. Shared token embedding and shared positional encoding are used since the encoder and decoder share the same amino acid tokens in the AE model.
To pretrain the AE for the proposed MMDesign framework solely based on the sequence data from the CATH training set (excluding the validation set and test set), we employ a sequence-to-sequence recovery task as a pseudo-translation process in natural language, as shown in Figure 2. This enables the AE to learn contextual semantic knowledge. And cross-entropy loss is used to fit the probability distribution of the generated sequences. Formally, we assume that the input protein sequence of the encoder is denoted as $S_{\text{in}}=\{s_{1},s_{2},\cdots,s_{n}\}$ with $n$ amino acids, and the generated protein sequence of the decoder is denoted as $S_{\text{out}}=\{s^{\prime}_{1},s^{\prime}_{2},\cdots,s^{\prime}_{n}\}$, which corresponds to the probability distribution denoted as $S_{\text{logits}}$. The reference native sequence is $S_{\text{native}}=S_{\text{in}}=\{s_{1},s_{2},\cdots,s_{n}\}$. The sequence cross-entropy loss is as follows:
$\mathcal{L}_{\textrm{seqCE}}=\textrm{CE}\big{(}\textrm{Softmax}(S_{\textrm{logits}}),S_{\textrm{native}}\big{)}.$ (1)
Since a relatively small training set of only 18k instances is used, the AE network could converge quickly to obtain better initialization parameters.
3.2 Protein Design Framework Training
3.2.1 MMDesign Training Pipeline
In the MMDesign framework, the pretrained structural module and the pretrained contextual module are sequentially concatenated, as illustrated in the lower section of Figure 2. The pretrained contextual module only preserves the encoder and decoder layers of the AE model in step 1 and discards the embedding layers, acting as the repository of prior language knowledge.
[…]
Firstly, the protein backbone $X$ is input to the pretrained GVPConv structural module to obtain the geometric structural features $Z_{\text{struc}}=\{z_{0},z_{1},\cdots,z_{n}\}$ with a dimension $d$ as:
$Z_{\text{struc}}=\text{GVP}(X)\in\mathbb{R}^{n\times d}.$ (2)
Then, the structural features $Z_{\text{struc}}$ are fed to the encoder-decoder layers of the pretrained contextual module sequentially. The features obtained by the encoder are referred to contextual features $Z_{\text{seq}}=\{z^{\prime}_{0},z^{\prime}_{1},\cdots,z^{\prime}_{n}\}$ as:
$Z_{\text{seq}}=Encoder(Z_{\text{struc}})\in\mathbb{R}^{n\times d}.$ (3)
Finally, the decoder decodes the contextual features to generate the output distribution $D_{\text{logits}}=\{o_{0},o_{1},\cdots,o_{n}\}$ of the generated sequence as:
$\begin{split}D_{\text{logits}}&=Decoder(Z_{\text{seq}})\in\mathbb{R}^{n\times M},\\ Y&=\text{argmax}\big{(}\text{Softmax}(D_{\text{logits}})\big{)}\in\mathbb{R}^{n},\end{split}$ (4)
where $M$=20 denotes 20 types of the most common amino acids.
# Models Perplexity Recovery (%)
All Short Single-chain Ts50 Ts500 All Short Single-chain Ts50 Ts500
1 *Natural frequencies [8] 17.97 18.12 18.03 - - 9.5 9.6 9.0 - -
SPIN [17] - - - - - - - - 30.3 30.3
SPIN2 [21] - 12.11 12.61 - - - - - 33.6 36.6
Structured Transformer [13] 6.85 8.54 9.03 5.60 5.16 36.4 28.3 27.6 42.40 44.66
Structured GNN [14] 6.55 8.31 8.88 5.40 4.98 37.3 28.4 28.1 43.89 45.69
*ESM-IF [8] 6.44 8.18 6.33 - - 38.3 31.3 38.5 - -
GCV [24] 6.05 7.09 7.49 5.09 4.72 37.64 32.62 31.10 47.02 47.74
*GVP-GNN-large [14] 6.17 7.68 6.12 - - 39.2 32.6 39.4 - -
GVP-GNN [14] 5.29 7.10 7.44 4.71 4.20 40.2 32.1 32.0 44.14 49.14
AlphaDesign [5] 6.30 7.32 7.63 5.25 4.93 41.31 34.16 32.66 48.36 49.23
ProteinMPNN [4] 4.61 6.21 6.68 3.93 3.53 45.96 36.35 34.43 54.43 58.08
PiFold [6] 4.55 6.04 6.31 3.86 3.44 51.66 39.84 38.53 58.72 60.42
2 ESM-IF [8] 3.99 6.30 6.29 3.43 3.34 52.51 34.74 34.25 56.66 60.85
3 Ours (MMDesign) 3.86 5.32 5.43 2.89 3.20 54.88 39.89 39.24 59.87 61.06
Table 1: Perplexity and recovery scores for MMDesign and other baselines on the CATH, Ts50, and Ts500 datasets. The best results are bolded, and the second best underlined. Group 1 & Group 3 are trained on a small-scale CATH training set only, while Group 2 is trained on a big-scale CATH+AlaphaDB training set. The Short and Single-chain are subsets of CATH test set ALL. *: evaluated on CATH 4.3 test set; Others: evaluated on CATH 4.2 test set. (Due to version differences, several baselines used CATH 4.3. However, both versions are sampled in CATH, so the data distribution and conclusions are similar.)
3.2.2 Generation Objective
The cross entropy loss is commonly used in protein design tasks to measure the discrepancy between the probability distribution of the generated amino acid sequences and the reference native sequences. By minimizing the cross entropy, we can obtain an approximate distribution that better matches the target distribution. In our experiments, we observed that emphasizing the salient features of the output distribution can further improve the performance of our model. To achieve this, we propose an exponential cross entropy (expCE) that accentuates the spiking effect of the output distribution. The expCE is calculated as follows:
$\mathcal{L}_{\text{expCE}}=\exp\left(\sum_{b=0}^{B}\text{CE}(\text{Softmax}(D_{\text{logits}}^{(b)}),\hat{Y}^{(b)})\right),$ (5)
where $exp$ denotes the exponential calculation, and $b$ denotes the $b$-th element in a batch of $B$ elements.
3.2.3 Cross-modal Alignment
Multi-modality research has shown that explicitly enforcing consistency between different modalities can improve cross-modal frameworks [35, 20, 7, 30, 31, 33]. In this work, we view protein design as a cross-modal task, where the structural features $Z_{\text{struc}}$ obtained from the GVPConv module and the contextual features $Z_{\text{seq}}$ obtained from the encoder of the contextual module are considered as two different modalities. To enforce alignment constraints across modalities, we propose a cross-layer cross-modal alignment constraint (CAC) loss, as shown in Figure 2. The CAC loss is implemented as a knowledge distillation loss (KL-divergence here), where the entire deep contextual features and the structural features are treated as teacher and student models, respectively (the teacher model stops back-propagating gradients). A high temperature $T=8$ is adopted to “soften” the probability distribution from potential spike responses[34]. The distillation process is expressed as:
$\mathcal{L}_{\text{CAC}}=\text{KL}(\text{Softmax}(\frac{Z_{\text{struc}}}{T}),\text{Softmax}(\frac{Z_{\text{seq}}}{T})).$ (6)
Overall, to align the generated sequences’ probability distribution with the reference distribution, we use $\mathcal{L}_{\text{expCE}}$ to fit the probability distribution. To achieve cross-layer cross-modal alignment, we use the CAC loss to align the multi-modality features by providing long-range supervision to the structure extractor. The final objective function of MMDesign training is composed of the primary $\mathcal{L}_{\text{expCE}}$ loss and the auxiliary $\mathcal{L}_{\text{CAC}}$ loss.
Figure 3: Normalized statistics of the number of amino acid residue types corresponding to the ground-truth sequences (base) and generated sequences of protein design models on CATH test set. The KL value indicates the KL divergence of the generated sequence distribution with respect to the ground-truth sequence distribution.
4 Experiments
4.1 Settings
Datasets. The CATH 4.2 [13] dataset is a curated version of CATH [22], in which all available structures with 40% non-redundancy are partitioned by their CATH classification(class, architecture, topology/fold, homologous super-family). The training, validation, and testing splits contain 18204, 608, and 1120 structure-sequence pairs, respectively. The training split is used for training our MMDesign, and the validation and testing splits are for evaluation purposes. CATH specifically refers to CATH 4.2 if not otherwise specified. We also report results on Ts50 & Ts500 [17] to assess the generalization. No canonical training and validation sets exist for Ts50 or Ts500.
Evaluation Metrics. Perplexity and sequence recovery on held-out native sequences are two commonly adopted metrics for protein design tasks. Perplexity measures the inverse likelihood of native sequences in the predicted sequence distribution (lower perplexity for higher likelihood). And sequence recovery measures the accuracy with which the sampled sequence matches the native sequence at each site. To maximize sequence recovery, the predicted sequences are sampled with low temperature $61e-$ 6 from the model.
Implementation Details. During training, we employed the SGD optimizer with a batch size of 5 and a learning rate of 1e-3 to optimize the protein design model. The GVPConv module consists of 4 layers with a dropout of 0.1, and the top-k neighbors are set to 30. The node hidden dimensions of scalars and vectors are 1024 and 256, respectively. For the contextual Transformer module, both the encoder and decoder have 8 layers of 8-head multi-head self-attention blocks, with an embedding dimension of 512 and an attention dropout of 0.1. And 1 NVIDIA A100 80GB GPU is used.
Baselines. Table 1 lists typical and representative works, most of which are open source and easy to re-implement. Works in Group 1 and Group 3 are only trained on a small-scale CATH training set (approximately 18k training pairs). Earlier proposed baselines such as Structured Transformer[13] and GVP-GNN [14] are still competitive and influential with relatively good performances and lightweight advantages. ProteinMPNN [4] by Baker’s group is renowned for improving performance and speed and their impressive biological validation experiments. PiFold[6] follows closely behind with further enhancements in efficiency and effectiveness. ESM-IF[8] by Meta is singled out in Group 2 as a unique open-source representative with strong data augmentation (trained on approximately 12M training pairs from AlphaFoldDB[26]). In addition, our MMDesign’s architectures are similar to ESM-IF, thus ESM-IF is chosen as a baseline.
4.2 Main Results
In line with previous work[13, 14, 8, 6], we start with evaluation in the standard setting of sequence design given all backbone coordinates.
Evaluation on CATH Test Set. Table 1 reports the perplexity (lower is better) and recovery (higher is better) scores on the generic dataset CATH test set (“ALL”). MMDesign achieves the best performance with a perplexity of 3.86 and a recovery of 54.88%, outperforming the state-of-the-art ProteinMPNN, PiFold (Group 1), and ESM-IF (Group 2) by a significant margin.
Evaluation on Single/Short Chains. Following [13], we also evaluate the performance of the models on the subsets of the CATH test set, i.e., “Short” dataset (protein sequence length $\leq$ 100 residues) and “Single-chain” dataset (single-chain proteins recorded in the Protein Data Bank), for evaluation, as shown in Table 1. MMDesign achieves excellent performances on both “Single-chain” and “Short” datasets (although the recovery score of “Single-chain” is slightly lower than GVP-GNN-large on CATH 4.3). In particular, in terms of perplexity, MMDesign (Short: 5.32; Single-chain: 5.43) significantly outperforms PiFold (Short: 6.04; Single-chain: 6.31) and ProteinMPNN (Short: 6.21; Single-chain: 6.68).
Evaluation on Generalization. To fully compare the generalizability of the different models on the out-of-domain datasets, we report the results on the TS50 and TS500 datasets in Table 1. MMDesign still achieves consistent improvements for TS50 and TS500, where the recovery score in Ts50 is close to 60%, and the score in TS500 breaks 61%.
Overall, ESM-IF performs better in generalization on both generic and out-of-domain datasets, while PiFold and ProteinMPNN perform better on short and single-chain datasets. However, MMDesign outperforms all baseline models across a variety of test sets, including generic datasets, out-of-domain datasets, and short and single-chain datasets with balanced performances, showing the strong generalization and robustness of MMDesign. Notably, our results also demonstrate that a pretrained prior knowledge migration-based strategy with small data training can outperform the large data training strategy such as ESM-IF.
# PSM PCM Perplexity Recovery(%)
Dev Test Dev Test
1 6.93 6.76 35.40 36.86
2 6.30 6.22 39.94 41.81
3 4.75 4.52 47.05 49.41
4 4.23 3.86 52.67 54.88
Table 2: Ablation study of various pretrained modules based on the protein design framework on CATH validation set and test set (PSM: pretrained structural module; PCM: pretrained contextual module). The best results are bolded.
4.3 Pretrained Module Ablation
To illustrate the significance of the pretrained modules in our MMDesign framework, we conduct an ablation study on the CATH validation set and test set to evaluate the impact of the pretrained structural module (PSM) and the pretrained contextual module (PCM). Overall, both pretrained modules have enhanced effects on the performance, as shown in Table 2.
Comparing #1 vs. #3 and #1 vs. #2, we observe that the pretrained structural features contribute more to the performance. This observation is consistent with our motivation for introducing prior structural knowledge. Moreover, it is likely that the PCM module, as a downstream module of PSM, is directly influenced by the representation capability of PSM.
Surprisingly, comparing only #1 vs. #2, we find that performances of MMDesign are improved considerably even without using initialization parameters of the pretrained structural module, i.e., by using only the PCM. Interestingly, this is a completely new observation that has not been extensively explored in previous protein design studies. These findings suggest that the pretrained contextual module has learned rich knowledge of the protein language to improve protein design tasks. Comparing #3 with #4 further confirms this observation.
Figure 4: The designable analysis of different models based on 3D protein structural folding. TM score and GDT score are used to reflect the biologically reasonable performance.
4.4 Generated Sequence Distribution
As shown in Figure 3, we counted both the generated protein residues and raw protein residues according to amino acid residue categories. We generated approximately 181k amino acid residues for each model based on 1,120 protein sequences from the CATH test set. Representative protein design models are chosen for comparison with our MMDesign model. To facilitate comparison, we normalized the distribution using the maximum-minimum normalization process. Compared to the ground-truth distribution (base) derived from raw reference sequences in Figure 3 (green), our MMDesign and PiFold produce distributions that are visually closer to the base. To quantitatively characterize the similarity of the generated sequences to the base distribution, KL-divergence is employed. The results are consistent with the visual observations, where our MMDesign and PiFold have lower KL-divergence values (around 0.03) than other baselines, indicating better similarity to the reference distribution. In contrast, GVP-GNN and Structured Transformer perform poorly, with higher KL-divergence values.
4.5 Designable Analysis
We assume that a biologically reasonable protein design should satisfy the condition that the generated protein sequence can be folded into a designable backbone structure and is physically realizable. Otherwise, the generated protein sequence is considered useless. Building upon previous biological and computational approaches [25, 16, 28], we propose to use ESMFold [18] to fold the generated sequences into their corresponding 3D structures. TM score and GDT score [29] are used to evaluate the structural similarity between the predicted backbone structures against the original native structures.
To our knowledge, we introduce this designable evaluation for the first time to explore the biological feasibility of existing protein design tools. PiFold is excluded from this analysis due to its discarding of coordinate missing nodes, resulting in discontinuous regions that make it difficult to align native structures for a fair comparison with other models. As shown in Figure 4, we first perform a folding procedure on the reference native protein sequences to generate 3D structures that demonstrate a high-precision folding level close to the experimentally validated level (“base” in Figure 4) with a 0.8068 TM score and a 0.7896 GDT score, which is a standard reference here.
We then use selected protein design models (MMDesign, ESM-IF, GVP-GNN, Structured Transformer) to generate sequences on the CATH test set, which are subsequently folded back into 3D structures via ESMFold. Evaluation of the predicted protein backbones reveals that the folding performances of the sequences generated by MMDesign (0.7920 TM score and 0.7755 GDT score) and ESM-IF (0.7916 TM score and 0.7737 GDT score) are significantly better than that of the comparative GVP-GNN and Structured Transformer, approaching the level of “base”, which further demonstrates that the high recovery level of sequences generated by MMDesign is biologically reasonable.
5 Conclusion
The proposed MMDesign represents a significant step forward in protein design by introducing a novel multi-modality transfer learning approach and an explicit consistency constraint. Trained only on the small dataset CATH, MMDesign improves significantly over baselines across various benchmarks. We also systematically present internal and external analysis techniques based on biological reasonable perspectives and data distribution statistics, which greatly improve the interpretability of models and reveal more about the nature and laws of protein design.
In future work, it would be interesting to explore the use of large language models to guide the learning of the alignment between structural and contextual modalities in an end-to-end manner, which could simplify the training pipeline and potentially further improve the performance of MMDesign. Additionally, it would be valuable to evaluate MMDesign on larger and more diverse datasets to further validate its effectiveness and generalizability.